Use the pumping lemma for context-free languages to prove that L3 is not a CFL.

L3 = { w: w e{a,b,c}* and na(w) < nh(w) < nc(w) }.
Answer #1

Let us assume that the string z = anbn+1cn+2 belongs to L.

Let us assume that L is context-free.

Now, let us assume that u = ap, v = an-p, w = bq, x = bn+1-q, y = cn+2

Now, in order for L to be CFL, the string uviwxiy should belong to L too.

uviwxiy = a(2ni - pi + p) b(2ni - qi + q + 1) c(n + 2)

Clearly, for larger values of i, we cannot guarantee that

(2ni - pi + p) < (n+2) or (2ni - qi + q + 1) <= (n + 2).

Hence, L is not CFL.